OpenStudy (anonymous):

What is the solution to the equation 2^ (1 + 2x) = 3 ? x ≈ –0.292 x ≈ –0.185 x ≈ 0.185 x ≈ 0.292 I got the last one as my answer, am I right? What is the solution to the equation 2^ (1 + 2x) = 3 ? x ≈ –0.292 x ≈ –0.185 x ≈ 0.185 x ≈ 0.292 I got the last one as my answer, am I right? @Mathematics

OpenStudy (across):

All your questions scream "logarithms!" It would be a good idea to review some of their properties. :)

myininaya (myininaya):

\[\log_2(2^{1+2x})=\log_2(3)\] \[(1+2x) \log_2(2)=\log_2(3)\] \[(1+2x)(1)=\log_2(3)\] \[1+2x=\log_2(3)\] \[2x=\log_2(3)-1\] \[x=\frac{\log_2(3)-1}{2}=\frac{\log_2(3)}{2}-\frac{1}{2}=\frac{1}{2} \cdot \frac{\ln(3)}{\ln(2)} -\frac{1}{2}\]

myininaya (myininaya):

yes last one

OpenStudy (anonymous):

Thanks :)